Solve theequation A = 2 (a + b -5), for b.

Answer by mananth(16949)

A = 2 (a + b -5), for b.
A=2a+2b-10
..
add 10-2a to both sides of the equation
A+10-2a= 2a+2b-10+10-2a
A+10-2a=2b
divide by 2
(A+10-2a)/2 = b.
IF capital A is a typing error
then the answer is
(10-a)/2